# Probing Mechanical Reasoning in Large Vision Language Models

### v2

@@ 1 Introduction
Humans are uniquely capable of working with complex mechanical systems, ranging from routine tasks, such as assembling furniture, to large-scale civil endeavors, such as designing architectural structures and developing advanced technologies [Harari, 2014]. These capabilities are underpinned by the cognitive ability to reason about the relationships and interactions of physical objects, an ability known as mechanical reasoning [Clark, 2010, Harman, 2011, Vaesen, 2012]. While there is evidence that some animal species demonstrate limited mechanical reasoning, such as understanding basic tool use or object interactions [Shumaker et al., 2011], the human capacity for mechanical reasoning is unparalleled in its flexibility, sophistication, and creative potential. This distinction allows humans to innovate, solve complex problems, and adapt tools and systems to a wide range of environments and challenges [Allen et al., 2020, Allen, 2021]. As such, mechanical reasoning is a cornerstone of human intelligence, driving technological and cultural progress throughout history. Implementing mechanical reasoning is therefore a vital step toward developing artificial intelligence systems capable of achieving human-level performance in real-life scenarios. Given the rapid advancements in large language models (LLMs), particularly their variants that support visual operations, i.e. vision language models (VLMs), mechanical reasoning appears to be a critical area of assessment to evaluate their current reasoning capabilities and identify potential limitations.
@@
Although mechanical reasoning is a high-level cognitive ability that does not emerge until late childhood, it has been found to rely heavily on more foundational cognitive strategies [Hegarty et al., 1988, Kim and Spelke, 1999, Allen, 2021]. In particular, decades of research in cognitive science suggest that mental simulation, the process of constructing and operating mental models of the world to guide reasoning, is critical for many aspects of mechanical reasoning [Hegarty, 2004]. For example, studies employing methods such as interviews and eye-tracking have revealed that individuals dynamically construct spatial representations of gear and pulley systems to infer their properties as they mentally simulate stages of motion [Hegarty, 1992, Lehrer and Schauble, 1998, Kubricht et al., 2017, Rozenblit et al., 2002]. On the other hand, while it remains an ongoing debate, several studies have indicated that the ability to construct mental models of the world to guide reasoning, particularly in the context of spatial dimensions, remains significantly underdeveloped even in state-of-the-art LLMs [Hao et al., 2023, Mitchell and Krakauer, 2023, Goddu et al., 2024, Gao et al., 2024]. Since mental simulations necessitate this model-based reasoning capacity on the visuospatial domain, a thorough evaluation of the mechanical reasoning capabilities of VLMs could therefore provide valuable insights into this debate given the critical role of mental simulation in solving mechanical reasoning tasks.
@@ 2.1 Cognitive Experiments
The understanding of stability is essential for reasoning about states of equilibrium within physical systems, such as predicting whether a stack of blocks will remain upright or collapse [McCloskey, 1983a]. In our experiments, models are presented with images of objects like stools with varying base widths or angles of inclination. The task involves selecting the most stable configuration. Stability reasoning encompasses factors such as the center of gravity, base area, and force distribution, making it a critical baseline for evaluating mechanical reasoning.
@@
Pulley systems are widely used in cognitive psychology and physics education to study how individuals reason about force and motion [Hegarty and Sims, 1994], requiring an understanding of force distribution and machinery functions. For example, a simple task might involve determining which pulley system requires less effort to lift a weight. By testing VLMs’ ability to distinguish between fixed and movable pulleys and to predict object movement, we evaluate their capacity to infer about dynamic relationships within real-life scenarios.
@@
Gear systems are deterministic mechanical setups governed by well-defined rules, such as adjacent gears rotating in opposite directions and gear ratios determining relative speeds. These properties make gears an ideal domain for testing logical and causal reasoning [Hegarty et al., 1988]. Tasks in this category involve analyzing diagrams of connected gears to predict their rotational direction and speed.
@@
The leverage principle illustrates the relationship between force, distance, and torque. Balance-scale experiments in cognitive psychology have shown how humans progressively develop an understanding of leverage through iterative learning and application [Peirce, 1992]. Tasks in this category include determining how shifting weights on a seesaw or applying force to a lever affects balance.
@@
Inertia and motion are dynamic aspects of mechanical reasoning that require understanding how forces influence the movement of objects over time. These concepts are central to Newtonian mechanics and intuitive physics [McCloskey, 1983b]. Human cognition integrates spatial and temporal information to make predictions about motion and forces, as seen in studies of tool use and physical reasoning [Allen, 2021]. Our experiments include scenarios such as predicting the trajectory of an object on a moving cart or identifying the kinetic energy distribution of a pendulum. These tasks thereby probe VLMs to integrate information about multiple aspects of the physical world.
@@
Fluid mechanics involves understanding the behavior of liquids under various conditions, such as flow, external pressures, and volume changes. Although gounded in the intuitive understanding of fluid dynamics emerged very early in humans’ cognitive development [Hespos et al., 2016], reasoning about liquid behaviors in mechanical systems require simultaneous consideration of geometry, force, and dynamics. These tasks represent a highly important dimension of mechanical reasoning concerning specifically about liquid motions as opposed to solid objects alone.
@@ 2.2 Model Selection and Experiment
    Open-source VLMs with Multi-Image Reasoning: Includes models with different sizes and other variants such as CogVLM Series[Hong et al., 2024], Qwen series(Qwen-VL [Bai et al., 2023], Qwen-2 [Wang et al., 2024]), and Blip2 [Li et al., 2023], LLaVA-Next [Liu et al., 2024] , which are capable of reasoning over interleaved multiple images and texts.
@@
    Closed-source VLMs with Multi-Image Reasoning: Includes proprietary models such as GPT series [OpenAI, ] ( GPT-4v, GPT-4-turbo, GPT-4o-mini), Gemini Series [Gemini, ], and Claude Series [claude, ]. These models also support reasoning across interleaved images and texts,
@@
    Open-source VLMs with single-Image Reasoning: Includes models designed to process a single image alongside continuous text. InstructBlip Series [Dai et al., 2023], LLaVA Series [Liu et al., 2023a] [Liu et al., 2023b]
In total, we aligned 26 models for evaluation. In order to analyze the reasoning abilities of VLMs, we asked the models to explain their answers after they have given the answers by adding "please provide an explanation" in the prompt.
@@ 2.3 Human Baseline
In addition to answering the experimental questions, participants completed a test of visuospatial fluid intelligence using Raven’s Advanced Progressive Matrices (APM), a widely used non-verbal assessment tool for evaluating fluid intelligence [Raven et al., 2000]. The APM results were collected to explore correlations between participants’ general visuospatial reasoning abilities and their performance on the experimental tasks.
@@ 3.1 General Results
Figure 4: Comparison of Tukey HSD P-values for Model and Human Performance Across Different Dimensions. A. The ANOVA results for model performance reveal a significant difference across the six dimensions (F-statistic = 11.8111, p < 0.0001), indicating that the models exhibit varying accuracy depending on the task. Tukey HSD tests show that significant differences exist between several pairs of dimensions, with System Stability, Inertia and Motion, and Fluid Systems exhibiting large disparities in accuracy. Leverage Principle and Gear Systems show less pronounced differences. B. Human performance across the six dimensions shows no significant difference (F-statistic = 1.6809, p = 0.1430). Tukey HSD tests for human performance indicate that none of the comparisons between different dimensions reach statistical significance, suggesting a more uniform performance across dimensions.
@@ 3.3 Relationship Between Model Performance and Model Size
A widely held belief in the machine learning community is that an increase in a model’s scale, measured by the number of parameters, leads to systematic improvements in its reasoning abilities [Sutton, 2019, Kaplan et al., 2020], a concept known as the scaling law. However, this assumption is an empirical observation without theoretical proof. To evaluate whether the scaling law applies to mechanical reasoning, we further examined the relationship between model performance on mechanical reasoning tasks and model size, as measured by the number of parameters (Figure 5). For overall accuracy across task, regression analysis yielded a formula $y=0.0821x+0.4318$ (p = 0.0053, R² = 0.2917). This trend was particularly evident in models with parameter counts exceeding 10 billion, such as GPT-4o High and GPT4V-TURBO-HIGH, which consistently outperformed smaller models. In most of the dimensions (System Stability, Pulley Systems, Leverage Principle, Inertia and Motion, Fluid Mechanics), accuracy increases with larger model sizes. However, Gear Systems and Fluid Systems show very weak relationships, with slopes and R-squared values close to zero (Gear Systems: $y=0.0010x+0.5547$, p = 0.981628, R² = 0.000023; Fluid Systems: $y=0.0169x+0.3695$, p = 0.787137, R² = 0.003097). These findings suggest that while model size often correlates with improved performance in many mechanical reasoning domains, some domains may not benefit significantly from increased model scale. This indicates a potential distinction between the underlying mechanisms across different domains of mechanical reasoning.
@@ 4 Discussions
Notably, VLMs’ performance on two dimensions—Gear Systems and Fluid Mechanics—remains close to chance and does not improve with increased model size, serving as a counterexample to the scaling law hypothesis in machine learning [Sutton, 2019, Kaplan et al., 2020]. Such a trend indicate that the underlying architecture of these models is yet to be able to support certain mechanisms that are required for a system to have mechanical reasoning abilities regarding Gear System and Fluid Mechanics. Similar trends have previously been reported in VLMs’ performance on level-2 perspective-taking tasks, in which the models are asked to tell how another agent would see the same spatial arrangement in a perspective different from their own [Gao et al., 2024, Piaget and Inhelder, 1969, Moll and Meltzoff, 2011]. Interestingly, like mechanical reasoning, said ability is also known to be supported by mental simulation in humans [Hegarty, 2004, Zhao et al., 2016]. Together, these findings suggest that VLMs may lack the ability to perform model-based reasoning, possibly highlighting a fundamental limitation in the architecture of the foundational models used in current VLMs.
@@
Figure 6: The Relationship Between Human Performance on MechBench and Human Fluid Intelligence. We further validate our MechBench using correlations with human fluid intelligence, measured via Raven’s Advanced Progressive Matrices [Raven et al., 2000]. Interestingly, System Ability seems to be an innate cognitive ability that’s orthogonal to human fluid intelligence.
@@
There is however an important concern with this interpretation: if the limitations of VLMs in mechanical reasoning are indeed due to their inability to perform model-based reasoning, why do they not also under-perform in domains like Pulley Systems, which are also known to rely on such reasoning? Two possible explanations could account for this discrepancy. Foremost, the role of mental simulations in mechanical reasoning is known to be mediated by low-level cognitive abilities, such as intuitions about object relations and motions—often referred to as intuitive physics [McCloskey, 1983a, Kubricht et al., 2017]. Intuitive physics involves understanding how objects and substances in the environment are subject to physical laws without relying on formal, abstract knowledge. While intuitive physics may draw on domain-general abilities like model-based reasoning, it also depends on the differential application of key physical principles, such as motion, gravity, and states of matter [Kaiser et al., 1986, Kubricht et al., 2017]. Different domains of mechanical reasoning are supported by an intuitive understanding of these physical principles. For instance, reasoning about fluid mechanics heavily relies on understanding liquid properties, which differentiates from that of solid objects [Kawabe et al., 2015]. It is possible that models can perform model-based reasoning at a lower level but fail to apply it effectively to certain physical domains, explaining their varied performance across mechanical reasoning tasks.
@@
On the other hand, it may just be that models have learned shortcuts that allow them to bypass model-based reasoning when tackling certain domains. Pulley Systems, for example, are heavily represented in online materials related to physics education. Given that Pulley Systems often involve a small number of components (e.g., fixed and moving pulleys), with limited variety in their arrangements, models may exploit spurious correlations between benchmark questions and similar examples already present in their training data. In other words, the performance of models on certain domains of mechanical reasoning may not reflect genuine, step-by-step reasoning but rather statistical correlations [Bender et al., 2021, Bleeker et al., 2024]. This is supported by the observation that very large models, such as GPT-4o, perform particularly well on Pulley System tasks, while smaller models tend to perform much worse.
@@ 5 Conclusion
Acknowledgments and Disclosure of Funding
The project is not funded by any organizations.
@@ A.1 System Stability
VLMs performance in intuitively evaluating system stability is not ideal, however here we observe some interesting phenomena. Models excel at identifying objects in images, and they not only understand what the objects are, the models also successfully recognize their mechanical states as well. Also, the models could effectively connect mechanical descriptions with the corresponding mechanical scenarios (Figure 7). In Experiments 7B and 7F, the models could recognize the two chairs and the two bottles, and their mechanical situations; however, the models still failed to provide the correct answers in the experiments. For instance, in Experiment 7B, the model explains, "the stool on the right is more likely to tip over when an active child sits on it because its legs are splayed at an angle, making the base wider and potentially more stable under normal circumstances but also adding a tipping hazard due to the non-vertical configuration of the legs". The model correctly notices that it’s the angle of the leg that matters for the stability of the system. However, it reasons completely the opposite way to correct answers. When the leg’s angle wider, it’s actually more stable. It’s a very intuitive physical problem for humans but the models fail, even though they still demonstrate step-by-step reasoning abilities in this case. In Experiment 7F, the model is correct that the bottle on the bottom "has a larger area of contact with the surface, creating more friction, which helps prevent rolling". However, the model fails to realize one bottle is standing, and one bottle is rolling, and sliding friction and rolling friction are completely different. In contrast, humans can easily solve these problems intuitively.

### v3

@@ 1 Introduction
Humans are uniquely capable of working with complex mechanical systems, ranging from routine tasks, such as assembling furniture, to large-scale civil endeavors, such as designing architectural structures and developing advanced technologies (Harari, 2014). These capabilities are underpinned by the cognitive ability to reason about the relationships and interactions of physical objects, an ability known as mechanical reasoning (Clark, 2010; Harman, 2011; Vaesen, 2012). While there is evidence that some animal species demonstrate limited mechanical reasoning, such as understanding basic tool use or object interactions (Shumaker et al., 2011), the human capacity for mechanical reasoning is unparalleled in its flexibility, sophistication, and creative potential. This distinction allows humans to innovate, solve complex problems, and adapt tools and systems to a wide range of environments and challenges (Allen et al., 2020; Allen, 2021). As such, mechanical reasoning is a cornerstone of human intelligence, driving technological and cultural progress throughout history. Implementing mechanical reasoning is therefore a vital step toward developing artificial intelligence systems capable of achieving human-level performance in real-life scenarios. Given the rapid advancements in large language models (LLMs), particularly their variants that support visual operations, i.e. vision language models (VLMs), mechanical reasoning appears to be a critical area of assessment to evaluate their current reasoning capabilities and identify potential limitations.
@@
Although mechanical reasoning is a high-level cognitive ability that does not emerge until late childhood, it has been found to rely heavily on more foundational cognitive strategies (Hegarty et al., 1988; Kim & Spelke, 1999; Allen, 2021). In particular, decades of research in cognitive science suggest that mental simulation, the process of constructing and operating mental models of the world to guide reasoning, is critical for many aspects of mechanical reasoning (Hegarty, 2004). For example, studies employing methods such as interviews and eye-tracking have revealed that individuals dynamically construct spatial representations of gear and pulley systems to infer their properties as they mentally simulate stages of motion (Hegarty, 1992; Lehrer & Schauble, 1998; Kubricht et al., 2017; Rozenblit et al., 2002). On the other hand, while it remains an ongoing debate, several studies have indicated that the ability to construct mental models of the world to guide reasoning, particularly in the context of spatial dimensions, remains significantly underdeveloped even in state-of-the-art LLMs (Hao et al., 2023; Mitchell & Krakauer, 2023; Goddu et al., 2024; Gao et al., 2024). Since mental simulations necessitate this model-based reasoning capacity on the visuospatial domain, a thorough evaluation of the mechanical reasoning capabilities of VLMs could therefore provide valuable insights into this debate given the critical role of mental simulation in solving mechanical reasoning tasks.
@@ 2.1 Cognitive Experiments
The understanding of stability is essential for reasoning about states of equilibrium within physical systems, such as predicting whether a stack of blocks will remain upright or collapse (McCloskey, 1983a). In our experiments, models are presented with images of objects like stools with varying base widths or angles of inclination. The task involves selecting the most stable configuration. Stability reasoning encompasses factors such as the center of gravity, base area, and force distribution, making it a critical baseline for evaluating mechanical reasoning.
@@
Pulley systems are widely used in cognitive psychology and physics education to study how individuals reason about force and motion (Hegarty & Sims, 1994), requiring an understanding of force distribution and machinery functions. For example, a simple task might involve determining which pulley system requires less effort to lift a weight. By testing VLMs’ ability to distinguish between fixed and movable pulleys and to predict object movement, we evaluate their capacity to infer about dynamic relationships within real-life scenarios.
@@
Gear systems are deterministic mechanical setups governed by well-defined rules, such as adjacent gears rotating in opposite directions and gear ratios determining relative speeds. These properties make gears an ideal domain for testing logical and causal reasoning (Hegarty et al., 1988). Tasks in this category involve analyzing diagrams of connected gears to predict their rotational direction and speed.
@@
The leverage principle illustrates the relationship between force, distance, and torque. Balance-scale experiments in cognitive psychology have shown how humans progressively develop an understanding of leverage through iterative learning and application (Peirce, 1992). Tasks in this category include determining how shifting weights on a seesaw or applying force to a lever affects balance.
@@
Inertia and motion are dynamic aspects of mechanical reasoning that require understanding how forces influence the movement of objects over time. These concepts are central to Newtonian mechanics and intuitive physics (McCloskey, 1983b). Human cognition integrates spatial and temporal information to make predictions about motion and forces, as seen in studies of tool use and physical reasoning (Allen, 2021). Our experiments include scenarios such as predicting the trajectory of an object on a moving cart or identifying the kinetic energy distribution of a pendulum. These tasks thereby probe VLMs to integrate information about multiple aspects of the physical world.
@@
Fluid mechanics involves understanding the behavior of liquids under various conditions, such as flow, external pressures, and volume changes. Although gounded in the intuitive understanding of fluid dynamics emerged very early in humans’ cognitive development (Hespos et al., 2016), reasoning about liquid behaviors in mechanical systems require simultaneous consideration of geometry, force, and dynamics. These tasks represent a highly important dimension of mechanical reasoning concerning specifically about liquid motions as opposed to solid objects alone.
@@ 2.2 Model Selection and Experiment
    Open-source VLMs with Multi-Image Reasoning: Includes models with different sizes and other variants such as CogVLM Series(Hong et al., 2024), Qwen series(Qwen-VL (Bai et al., 2023), Qwen-2 (Wang et al., 2024)), and Blip2 (Li et al., 2023), LLaVA-Next (Liu et al., 2024) , which are capable of reasoning over interleaved multiple images and texts.
@@
    Closed-source VLMs with Multi-Image Reasoning: Includes proprietary models such as GPT series (OpenAI, ) ( GPT-4v, GPT-4-turbo, GPT-4o-mini), Gemini Series (Gemini, ), and Claude Series (claude, ). These models also support reasoning across interleaved images and texts,
@@
    Open-source VLMs with single-Image Reasoning: Includes models designed to process a single image alongside continuous text. InstructBlip Series (Dai et al., 2023), LLaVA Series (Liu et al., 2023a) (Liu et al., 2023b)
In total, we aligned 26 models for evaluation. In order to analyze the reasoning abilities of VLMs, we asked the models to explain their answers after they have given the answers by adding ”please provide an explanation” in the prompt.
@@ 2.3 Human Baseline
In addition to answering the experimental questions, participants completed a test of visuospatial fluid intelligence using Raven’s Advanced Progressive Matrices (APM), a widely used non-verbal assessment tool for evaluating fluid intelligence (Raven et al., 2000). The APM results were collected to explore correlations between participants’ general visuospatial reasoning abilities and their performance on the experimental tasks.
@@ 3.1 General Results
Figure 4: Comparison of Tukey HSD P-values for Model and Human Performance Across Different Dimensions. A. The ANOVA results for model performance reveal a significant difference across the six dimensions (F-statistic = 11.8111, p ¡ 0.0001), indicating that the models exhibit varying accuracy depending on the task. Tukey HSD tests show that significant differences exist between several pairs of dimensions, with System Stability, Inertia and Motion, and Fluid Systems exhibiting large disparities in accuracy. Leverage Principle and Gear Systems show less pronounced differences. B. Human performance across the six dimensions shows no significant difference (F-statistic = 1.6809, p = 0.1430). Tukey HSD tests for human performance indicate that none of the comparisons between different dimensions reach statistical significance, suggesting a more uniform performance across dimensions.
@@ 3.3 Relationship Between Model Performance and Model Size
A widely held belief in the machine learning community is that an increase in a model’s scale, measured by the number of parameters, leads to systematic improvements in its reasoning abilities (Sutton, 2019; Kaplan et al., 2020), a concept known as the scaling law. However, this assumption is an empirical observation without theoretical proof. To evaluate whether the scaling law applies to mechanical reasoning, we further examined the relationship between model performance on mechanical reasoning tasks and model size, as measured by the number of parameters (Figure 5). For overall accuracy across task, regression analysis yielded a formula $y=0.0821x+0.4318$ (p = 0.0053, R² = 0.2917). This trend was particularly evident in models with parameter counts exceeding 10 billion, such as GPT-4o High and GPT4V-TURBO-HIGH, which consistently outperformed smaller models. In most of the dimensions (System Stability, Pulley Systems, Leverage Principle, Inertia and Motion, Fluid Mechanics), accuracy increases with larger model sizes. However, Gear Systems and Fluid Systems show very weak relationships, with slopes and R-squared values close to zero (Gear Systems: $y=0.0010x+0.5547$, p = 0.981628, R² = 0.000023; Fluid Systems: $y=0.0169x+0.3695$, p = 0.787137, R² = 0.003097). These findings suggest that while model size often correlates with improved performance in many mechanical reasoning domains, some domains may not benefit significantly from increased model scale. This indicates a potential distinction between the underlying mechanisms across different domains of mechanical reasoning.
@@ 4 Discussions
Notably, VLMs’ performance on two dimensions—Gear Systems and Fluid Mechanics—remains close to chance and does not improve with increased model size, serving as a counterexample to the scaling law hypothesis in machine learning (Sutton, 2019; Kaplan et al., 2020). Such a trend indicate that the underlying architecture of these models is yet to be able to support certain mechanisms that are required for a system to have mechanical reasoning abilities regarding Gear System and Fluid Mechanics. Similar trends have previously been reported in VLMs’ performance on level-2 perspective-taking tasks, in which the models are asked to tell how another agent would see the same spatial arrangement in a perspective different from their own (Gao et al., 2024; Piaget & Inhelder, 1969; Moll & Meltzoff, 2011). Interestingly, like mechanical reasoning, said ability is also known to be supported by mental simulation in humans (Hegarty, 2004; Zhao et al., 2016). Together, these findings suggest that VLMs may lack the ability to perform model-based reasoning, possibly highlighting a fundamental limitation in the architecture of the foundational models used in current VLMs.
@@
Figure 6: The Relationship Between Human Performance on MechBench and Human Fluid Intelligence. We further validate our MechBench using correlations with human fluid intelligence, measured via Raven’s Advanced Progressive Matrices (Raven et al., 2000). Interestingly, System Ability seems to be an innate cognitive ability that’s orthogonal to human fluid intelligence.
@@
There is however an important concern with this interpretation: if the limitations of VLMs in mechanical reasoning are indeed due to their inability to perform model-based reasoning, why do they not also under-perform in domains like Pulley Systems, which are also known to rely on such reasoning? Two possible explanations could account for this discrepancy. Foremost, the role of mental simulations in mechanical reasoning is known to be mediated by low-level cognitive abilities, such as intuitions about object relations and motions—often referred to as intuitive physics (McCloskey, 1983a; Kubricht et al., 2017). Intuitive physics involves understanding how objects and substances in the environment are subject to physical laws without relying on formal, abstract knowledge. While intuitive physics may draw on domain-general abilities like model-based reasoning, it also depends on the differential application of key physical principles, such as motion, gravity, and states of matter (Kaiser et al., 1986; Kubricht et al., 2017). Different domains of mechanical reasoning are supported by an intuitive understanding of these physical principles. For instance, reasoning about fluid mechanics heavily relies on understanding liquid properties, which differentiates from that of solid objects (Kawabe et al., 2015). It is possible that models can perform model-based reasoning at a lower level but fail to apply it effectively to certain physical domains, explaining their varied performance across mechanical reasoning tasks.
@@
On the other hand, it may just be that models have learned shortcuts that allow them to bypass model-based reasoning when tackling certain domains. Pulley Systems, for example, are heavily represented in online materials related to physics education. Given that Pulley Systems often involve a small number of components (e.g., fixed and moving pulleys), with limited variety in their arrangements, models may exploit spurious correlations between benchmark questions and similar examples already present in their training data. In other words, the performance of models on certain domains of mechanical reasoning may not reflect genuine, step-by-step reasoning but rather statistical correlations (Bender et al., 2021; Bleeker et al., 2024). This is supported by the observation that very large models, such as GPT-4o, perform particularly well on Pulley System tasks, while smaller models tend to perform much worse.
@@ A.1 System Stability
VLMs performance in intuitively evaluating system stability is not ideal, however here we observe some interesting phenomena. Models excel at identifying objects in images, and they not only understand what the objects are, the models also successfully recognize their mechanical states as well. Also, the models could effectively connect mechanical descriptions with the corresponding mechanical scenarios (Figure 7). In Experiments 7B and 7F, the models could recognize the two chairs and the two bottles, and their mechanical situations; however, the models still failed to provide the correct answers in the experiments. For instance, in Experiment 7B, the model explains, ”the stool on the right is more likely to tip over when an active child sits on it because its legs are splayed at an angle, making the base wider and potentially more stable under normal circumstances but also adding a tipping hazard due to the non-vertical configuration of the legs”. The model correctly notices that it’s the angle of the leg that matters for the stability of the system. However, it reasons completely the opposite way to correct answers. When the leg’s angle wider, it’s actually more stable. It’s a very intuitive physical problem for humans but the models fail, even though they still demonstrate step-by-step reasoning abilities in this case. In Experiment 7F, the model is correct that the bottle on the bottom ”has a larger area of contact with the surface, creating more friction, which helps prevent rolling”. However, the model fails to realize one bottle is standing, and one bottle is rolling, and sliding friction and rolling friction are completely different. In contrast, humans can easily solve these problems intuitively.